If P(A)=0.60 and the probability that both events(eventA and eventB) occur simultaneously is 0.30 then P(B/A) is equal to
Answer #1

#Given:

P(A)=0.60

P(A and B)=0.30

P(B/A) =P(A and B)/P(A)

=0.30/0.60

=0.5

#P(B/A)=0.5

#P(B/A) is equal to 0.5